OpenStudy (anonymous):

Please help! Evan opens a savings account with $5,000. He deposits $75 every month into the account that compounds annually and has a 0.85% interest rate. What will his account total be in 5 years?

OpenStudy (anonymous):

P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year P = 5000 r = 8.5 / 100 = 0.085 t = 5 A = ??? n = 1

OpenStudy (amistre64):

i dont see a payment option in that setup

OpenStudy (amistre64):

i would just run a setup of adding payments instead of subtracting them; like this Bn = Bo k^t + P(1-k^t)/(1-k) where k = 1+r

OpenStudy (amistre64):

might want to make P = 12(75) tho since to adjust for monthly payments

OpenStudy (amistre64):

M0 = 5000 M1 = 5000 + 75 M2 = 5000 + 75 + 75 M3 = 5000 + 3(75) M12 = 5000 + 12(75) ; then compound (5000 + 12(75))(1+.0085) M13 = (5000 + 12(75))(1+.0085) + 1(75) M14 = (5000 + 12(75))(1+.0085) + 2(75) M15 = (5000 + 12(75))(1+.0085) + 3(75) M24 = (5000 + 12(75))(1+.0085) + 12(75), and compound ((5000 + 12(75))(1+.0085) + 12(75))(1+.0085) let k = 1+.0085 for a clean up ((5000 + 12(75))(k) + 12(75))(k) (5000(k) + 12(75)(k) + 12(75)) (k) 5000(k)^2 + 12(75)(k) + 12(75)(k)^2 hmmm, maybe a geometric sum with (n+1) ?

OpenStudy (amistre64):

seems to be writing up like an annuity due structure

OpenStudy (amistre64):

or some hybrid monstrocity :)